Show that superposition principle does not hold for a non linear DE like; d²x/d²t + cx² = 0 where c is constant.

1
Expert's answer
2022-02-13T12:15:03-0500

Let us assume "x_1" and "x_2" both are solution of DE

"d\u00b2x\/d\u00b2t + cx\u00b2 = 0"

So

"d²x_1/d²t + cx_1^2= 0,\quad d²x_2/d²t + cx_2^2= 0"

Now, we consider a superposition of these solutions

"x=x_1+x_2"

We get

"d\u00b2(x_1+x_2)\/d\u00b2t + c(x_1+x_2)^2"

"=d²x_1/d²t + cx_1^2+d²x_2/d²t + cx_2^2+2cx_1x_2\\
=2cx_1x_2\neq0"

Thus, the superposition principle does not hold for a non linear DE.